# Strategic inputs: feature selection from game-theoretic perspective

### v1

@@ 1 Introduction
Non game-theoretic methods, such as Feature Importance (FI) (Breiman, 2001), Permutation Feature Importance (PFI) (Altmann et al., 2010), and Local Interpretable Model-Agnostic Explanations (LIME) (Ribeiro et al., 2016) are widely used in practice due to their computational efficiency and ease of implementation. However, these methods typically evaluate feature importance independently of other features, i.e., they cannot capture complex interactions between features.
Game-theoretic methods, particularly, those based on the Shapley value, provide a more comprehensive framework for feature importance evaluation. The SHAP method (SHapley Additive exPlanations) (Lundberg & Lee, 2017) provides a unified approach with various implementations including KernelSHAP, SamplingSHAP, etc. Recently developed ShapG method (Zhao et al., 2025a; b), inspired by Shapley value-based centrality measures on graphs (Zhao & Parilina, 2024), demonstrates improved computational efficiency compared to traditional Shapley-based methods. These game-theoretic approaches evaluate feature importance based on cooperative interactions among features, enabling more detailed understanding of feature contributions.
Despite their theoretical advantages, game-theoretic methods face significant computational challenges. The computational cost increases exponentially with the number of features when using the Shapley value to measure their importance. Both game-theoretic and non game-theoretic methods exhibit a linear relationship between data scale (number of observations) and computational cost for feature importance evaluation. Therefore, selecting a representative subset of observations can significantly reduce computational costs for feature importance evaluation. In the paper Axiotis et al. (2024), the authors proposed a clustering-based sensitivity sampling method to select diverse subsets of observations for efficient learning. In this work, we adopt a simpler clustering-based method that emphasizes balanced coverage of the data distribution, ensuring the sampled subset reflects both global and local diversity without over-emphasizing highly influential observations. The latter may lead to inaccurate feature importance estimators.
@@ 1. Step 1
    This can be solved by the $k$-means++ algorithm (see Lloyd (1982); Arthur & Vassilvitskii (2006)). By solving this minimization problem, we obtain the set of clusters $\{C_{1},\ldots,C_{k}\}$.
@@ 2.2 Feature selection
    Method 3. Global-model performance: Similar to equation 3, but using dataset-wide metrics: $R^{2}$, negative MAE, or negative RMSE for regression tasks; accuracy or F1 score for classification tasks (Zhao et al. (2025a)).
@@
The Shapley value (Shapley, 1951) is a widely used solution of this problem:
@@
The Centre of the Imputation Set (CIS) Value (Driessen & Funaki, 1991) offers a more computationally efficient alternative:
@@
To further improve computational efficiency, we adapt the graph-based sampling approach from the ShapG method (Zhao et al., 2025a). ShapG constructs a collaborator set $\mathcal{C}_{i}$ for each feature $i$, restricting feature interactions to within these sets during evaluation. When this set is not large, i.e. $|\mathcal{C}_{i}|<\ell$ (default $\ell=15$), the exact Shapley value can be computed; otherwise, Monte Carlo approximation is used. We integrate this approach into our pipeline without requiring model retraining, significantly reducing computational costs.
@@ 3 Experiments
In this section, we evaluate how our proposed feature selection pipeline affects the performance of deep-learning models on tabular data, focusing on MLP and TabM (Gorishniy et al., 2025), the current state-of-the-art on TabRed benchmarks (Rubachev et al., 2024).
@@ 3.1.1 Models
    MLP: A multilayer perceptron with fully connected layers [^1] We use the same architecture as in Gorishniy et al. (2025), the number of layers varies according to the dataset., ReLU activations, and task-specific output heads.
@@
    TabM: A parameter-efficient ensemble model for tabular data that trains multiple MLPs simultaneously while sharing most parameters between them, achieving state-of-the-art performance (Gorishniy et al., 2025).
@@
    MLP+PLE: MLP enhanced with PLE (Piecewise Linear Embedding, see Gorishniy et al. (2022)) for numerical features, which splits continuous values into segments and learns separate representations for each segment, enabling better modeling of non-linear relationships.
@@ 3.1.2 Configuration
Feature importance method ShapG
@@
We compare baseline models (all the features included) against models with feature selection (FS), reporting mean $\pm$ std of model performance and training time across 15 seeds.
@@ 3.2 Datasets
In total, we use 10 datasets from (Gorishniy et al., 2025).
Table 2: Dataset configuration from Gorishniy et al. (2025)
@@ 3.3 Implementation details
Training. We use the TabM benchmark suite with hyperparameters from Gorishniy et al. (2025), tuned via Optuna (Akiba et al., 2019).
Feature importance. We adapt ShapG (Zhao et al., 2025a; b) by moving evaluation from training to prediction stage, avoiding model retraining for each feature subset. Following Zhao et al. (2025b)’s improved strategy, we construct collaboration sets $\mathcal{C}_{i}$ through a graph-based approach: (1) rank features by their correlation with the target using cosine distance (1 - cosine similarity), (2) build a complete feature graph with edge weights based on pairwise cosine distance, (3) prune edges starting from the least correlated features (processed in ascending order of importance) while maintaining graph connectivity, targeting a graph density of $1.5\times$ the minimum connectivity ratio (capped at 1.0). This approach better captures feature interactions compared to the original method in Zhao et al. (2025a).
@@ 3.4 Results & analysis
Under the experimental setup described above, we evaluate performance of the proposed feature selection pipeline on 10 benchmark datasets. Our analysis reveals distinct patterns between classification and regression tasks, as shown in Figure 2. On the $OX$-axis of each graph, the four models (MLP, TabM, MLP+PLE, TabM+PLE) are presented. Ten subfigures correspond to ten datasets presented in Table 2. Accuracy (for classification tasks) or $R^{2}$ (for regression tasks) of the model is presented on the $OY$-axis. The purple (green) column corresponds to the baseline (top-80% feature) method.
Figure 2: Comparison of model performances across 10 datasets.
Classification tasks. Classification tasks demonstrate remarkable robustness to feature reduction, with accuracy degradation ranging from 0.5% to 8.7% and average 3% across all model-dataset combinations. We notice that the adult, otto, and covtype2 datasets maintain performance particularly well, with average accuracy drops below 2%. The most significant degradation occurs in the churn and higgs-small datasets. Our explanation of such a drop is that these tasks rely more significantly on the complete set of features.
Regression tasks. In contrast to classification, regression tasks exhibit substantially larger performance degradation under feature selection, with $R^{2}$ reductions ranging from 3.9% to 17.9%. The california and house datasets show the most pronounced sensitivity to feature reduction, with average $R^{2}$ drops exceeding 10%. The diamond dataset demonstrates relative stability, maintaining over 95% of its baseline $R^{2}$ performance. The potential reason for this sensitivity to reduction of features could be attributed to the bad quality of baseline models ($R^{2}$ is close to 1 for the diamond dataset, but it is significantly lower for other datasets).
Training time. Figure 3 presents the training time (presented on the $OY$-axis) comparison between baseline (purple columns) and top-80% feature (green columns) configurations. We observe that the training time for the top-80% feature configurations is not always lower than for the baseline ones. This can be explained by the fact that in our experiments, we did not set the limitation for epochs, we only have set the patience for early stopping. Therefore, the training time depends on the convergence of the model, in some cases, the model with fewer features converges faster, while in other cases, it takes longer time to converge.
Figure 3: Comparison of training time of different methods across 10 datasets.
Table 3: Summary of results: Baseline vs Top 80% Feature Selection
Model Classification (Accuracy) Regression ($R^{2}$) Overall Time
Baseline Top 80% $\Delta$% Baseline Top 80% $\Delta$% $\Delta$% $\Delta$%
MLP 0.8423 0.8157 -3.2 0.6327 0.5824 -7.9 -5.2 -7.2
MLP-PLE 0.8458 0.8257 -2.4 0.6448 0.5762 -10.6 -6.0 +1.3
TabM 0.8520 0.8254 -3.1 0.6507 0.5739 -11.8 -6.9 -19.9
TabM-PLE 0.8557 0.8264 -3.4 0.6568 0.6050 -7.9 -5.4 -23.2
Summary. Table 3 summarizes the results of our experiments. Feature selection successfully reduces the number of features while maintaining model performance comparable to using all features, with an average drop of 3% in classification accuracy, 9.6% in $R^{2}$ in regression task, and 5.9% overall. Additionally, training time is reduced by in average of 12.3%, though the effect varies with the model: MLP-PLE becomes slightly slower (+1.3%), MLP shows a modest speed-up (-7.2%), and both TabM and TabM-PLE achieve substantial reduction of time (-19.9% and -23.2%, respectively). These results allow to conclude that feature selection offers a favorable trade-off between predictive performance and computational efficiency, with particular benefits for more complex model architectures.
4 Conclusion and future work
@@
In this work, we proposed a feature selection pipeline based on game-theoretic feature importance evaluation. Our pipeline employs a k-means based diversity sampling procedure to select a representative subset of observations for feature importance evaluation, thereby reducing the influence of outliers and ensuring more robust feature ranking. We demonstrated the effectiveness of this approach on 10 benchmark datasets using state-of-the-art tabular deep learning models. The results show that integrating our feature selection pipeline into model training reduces computational costs by 12.3% on average while maintaining competitive performance, with accuracy degradation by 5.9% on average across all tasks.
@@
Our study has several limitations that open opportunities for future research. First, we evaluated only a single feature selection configuration (retaining the top 80% of features by importance) using the ShapG method, which showed effectiveness in recent feature importance evaluation studies Zhao et al. (2025a; b). Comprehensive comparisons with alternative feature selection approaches—including LIME, Permutation Feature Importance, and other game-theoretic methods with different characteristic functions—remain to be explored. We can also extend the proposed method by varying the number of features selected (the percentage of $q$-top features) for a fixed drop in the model performance. This extension is due to varying the number of significant features depending of the dataset and deep-learning model using as a basis.
@@
Additionally, extending this benchmark to industrial datasets, such as those described in Rubachev et al. (2024), would provide valuable insights on performance of real-world data analysis.
@@
The proposed diversity sampling method also shows promising results beyond feature importance evaluation. Preliminary experiments suggest its effectiveness for training data sampling, in particular, for domains with heterogeneous data distributions. While we have observed encouraging results on proprietary wireless communication datasets with TabM (Gorishniy et al., 2025) and FT-transformer (Gorishniy et al., 2021) models, we cannot present these findings due to confidentiality constraints. Future work should investigate this application systematically on publicly available datasets.
@@
These findings suggest that feature selection remains a valuable preprocessing step even for modern deep learning approaches to tabular data, offering a practical way to data analysts to balance computational efficiency with respect to predictive performance.

### v2

@@ 1 Introduction
Non game-theoretic methods, such as Feature Importance (FI) (breiman2001random), Permutation Feature Importance (PFI) (altmann2010permutation), and Local Interpretable Model-Agnostic Explanations (LIME) (ribeiro2016should) are widely used in practice due to their computational efficiency and ease of implementation. However, these methods typically evaluate feature importance independently of other features, i.e., they cannot capture complex interactions between features.
Game-theoretic methods, particularly, those based on the Shapley value, provide a more comprehensive framework for feature importance evaluation. The SHAP method (SHapley Additive exPlanations) (lundberg2017unified) provides a unified approach with various implementations including KernelSHAP, SamplingSHAP, etc. Recently developed ShapG method (zhao2024shapgnewfeatureimportance; zhao2025complete), inspired by Shapley value-based centrality measures on graphs (zhao2024centralitymeasuresopiniondynamics), demonstrates improved computational efficiency compared to traditional Shapley-based methods. A sparsity-based deterministic method for the Shapley value approximation presented in erofeeva2025novel introduces a novel compressed-sensing perspective on this problem. However, as the number of players (features) increases, the associated time and memory requirements grow rapidly, which may limit its applicability in high-dimensional settings. These game-theoretic approaches evaluate feature importance based on cooperative interactions among features, enabling more detailed understanding of feature contributions. A detailed description and benchmark results for each method can be found in (zhao2025shapley).
Despite their theoretical advantages, game-theoretic methods face significant computational challenges. The computational cost increases exponentially with the number of features when using the Shapley value to measure their importance. Both game-theoretic and non game-theoretic methods exhibit a linear relationship between data scale (number of observations) and computational cost for feature importance evaluation. Therefore, selecting a representative subset of observations can significantly reduce computational costs for feature importance evaluation. In the paper axiotis2024data, the authors proposed a clustering-based sensitivity sampling method to select diverse subsets of observations for efficient learning. In this work, we adopt a simpler clustering-based method that emphasizes balanced coverage of the data distribution, ensuring the sampled subset reflects both global and local diversity without over-emphasizing highly influential observations. The latter may lead to inaccurate feature importance estimators.
@@ 1. Step 1
    This can be solved by the $k$-means++ algorithm (see lloyd1982least; arthur2006k). By solving this minimization problem, we obtain the set of clusters $\{C_{1},\ldots,C_{k}\}$.
@@ 2.2 Feature selection
    Method 3. Global-model performance: Similar to equation 3, but using dataset-wide metrics: $R^{2}$, negative MAE, or negative RMSE for regression tasks; accuracy or F1 score for classification tasks (zhao2024shapgnewfeatureimportance).
@@
The Shapley value (shapley1951notes) is a widely used solution of this problem:
@@
The Centre of the Imputation Set (CIS) Value (driessen1991coincidence) offers a more computationally efficient alternative:
@@
To further improve computational efficiency, we adapt the graph-based sampling approach from the ShapG method (zhao2024shapgnewfeatureimportance). ShapG constructs a collaborator set $\mathcal{C}_{i}$ for each feature $i$, restricting feature interactions to within these sets during evaluation. When this set is not large, i.e. $|\mathcal{C}_{i}|<\ell$ (default $\ell=15$), the exact Shapley value can be computed; otherwise, Monte Carlo approximation is used. We integrate this approach into our pipeline without requiring model retraining, significantly reducing computational costs.
@@ 3 Experiments
In this section, we evaluate how our proposed feature selection pipeline affects the performance of deep-learning models on tabular data, focusing on MLP and TabM (gorishniy2024tabm), the current state-of-the-art on TabRed benchmarks (rubachev2024tabred).
@@ 3.1.1 Models
    MLP: A multilayer perceptron with fully connected layers [^1] We use the same architecture as in gorishniy2024tabm, the number of layers varies according to the dataset., ReLU activations, and task-specific output heads.
@@
    TabM: A parameter-efficient ensemble model for tabular data that trains multiple MLPs simultaneously while sharing most parameters between them, achieving state-of-the-art performance (gorishniy2024tabm).
@@
    MLP+PLE: MLP enhanced with PLE (Piecewise Linear Embedding, see gorishniy2022embeddings) for numerical features, which splits continuous values into segments and learns separate representations for each segment, enabling better modeling of non-linear relationships.
@@ 3.1.2 Configuration
Feature importance method ShapG and PFI
@@
We compare baseline models (all the features included) against models with feature selection (FS), reporting mean $\pm$ std of model performance and training time across 3 seeds.
@@ 3.2 Datasets
In total, we use 10 datasets from (gorishniy2024tabm).
Table 2: Dataset configuration from gorishniy2024tabm
@@ 3.3 Implementation details
Training. We use the TabM benchmark suite with hyperparameters from gorishniy2024tabm, tuned via Optuna (akiba2019optuna).
Feature importance. Since many datasets in our benchmark have more than 20 features, most existing feature-importance XAI methods are nearly infeasible in this setting. Thus, we focus on two methods for feature importance evaluation:
  1. 1.
    ShapG method: We adapt ShapG (zhao2024shapgnewfeatureimportance; zhao2025complete) by moving evaluation from training to prediction stage, avoiding model retraining for each feature subset. Following zhao2025complete’s improved strategy, we construct collaboration sets $\mathcal{C}_{i}$ through a graph-based approach: (1) rank features by their correlation with the target using cosine distance (1 - cosine similarity), (2) build a complete feature graph with edge weights based on pairwise cosine distance, (3) prune edges starting from the least correlated features (processed in ascending order of importance) while maintaining graph connectivity, targeting a graph density of $1.5\times$ the minimum connectivity ratio (capped at 1.0). This approach better captures feature interactions compared to the original method in zhao2024shapgnewfeatureimportance.
  2. 2.
    PFI method: We implement the Permutation Feature Importance (PFI) method (altmann2010permutation) by measuring the drop in a model performance when each feature values are randomly shuffled across samples. The algorithm shuffles each feature several times (5 by default), measures the effect after each shuffle, and averages the results to reduce randomness. Features giving larger drop in the performance when shuffled are deemed more important, providing a model-agnostic evaluation of feature importance without requiring model retraining.
@@ 3.4 Results & analysis
Under the experimental setup described above, we evaluate performance of the proposed feature selection pipeline on 10 benchmark datasets. Our analysis reveals distinct patterns between classification and regression tasks, as shown in Figure 2. On the $OX$-axis of each graph, the four models (MLP, TabM, MLP+PLE, TabM+PLE) are presented. Ten subfigures correspond to ten datasets presented in Table 2. Accuracy (for classification tasks) or $R^{2}$ (for regression tasks) of the model is presented on the $OY$-axis. The purple columns correspond to the baseline without feature selection, while the green and light-blue columns correspond to models using ShapG and Permutation feature selection (top 80% features), respectively.
Figure 2: Model performance ($R^{2}$ for regression task and accuracy for classification task) across 10 datasets.
Classification tasks. For classification tasks (datasets Adult, Churn Modelling, Covertype, Higgs Small, Otto Group Products), both feature-selection methods (ShapG and PFI) show consistently strong robustness to feature reduction. Across all datasets, both ShapG and Permutation selection of the top 80% most important features yield accuracy that remains close to the baseline models trained on all features (purple bars), with typical degradation of only a few percentage points and an average drop less than 4% across all model-dataset combinations. Notably, on datasets such as Adult, both ShapG and PFI maintain performance very close to the baseline. In some cases (datasets Covertype and Otto Group Products), ShapG significantly outperforms PFI in terms of average accuracy, whereas on the Higgs Small dataset, PFI substantially outperforms ShapG. For the Churn Modelling dataset, both methods exhibit comparable performance.
Regression tasks. In regression tasks (datasets Black Friday, California Housing, House 16H, Diamond, Microsoft), feature reduction affects the two methods very differently: ShapG shows substantial sensitivity, while the PFI method maintains stable performance. Across all regression datasets, ShapG exhibits substantial degradation in $R^{2}$, often exceeding 10%, that is observed in datasets California Housing and House 16H. In contrast, the PFI method remains much more stable: its performance drops are consistently small, typically below 6%, and in some cases even slightly surpass or remain close to the baseline (e.g., datasets California Housing and Diamond). This suggests that for regression tasks, the PFI method provides a more reliable feature importance evaluation compared to ShapG.
Training time. As shown in Figure 3, the impact of feature selection on training time varies across models and datasets. Notably, the 80% of feature configurations of both ShapG and PFI do not always lead to smaller training time compared with the baseline. This arises because the number of training epochs is not fixed in our setting; instead, we rely on early stopping with a patience parameter. Consequently, the actual training time is determined by the convergence dynamics of each model. In some cases, reducing the feature set accelerates convergence, whereas in others it may slow it down, leading to longer training time.
Figure 3: Training time for different methods across 10 datasets.
Table 3: Performance across different feature selection methods
XAI Method Model Classification Regression
Accuracy Time (s) $R^{2}$ Time (s)
Baseline MLP 0.8426 22.1 0.6329 16.9
MLP-PLE 0.8466 27.5 0.6443 24.7
TabM 0.8520 326.6 0.6510 277.1
TabM-PLE 0.8554 345.7 0.6561 328.8
ShapG MLP 0.8182 (-3.1%) 22.2 (-6.6%) 0.5865 (-7.1%) 13.9 (-8.4%)
MLP-PLE 0.8226 (-3.0%) 26.3 (+4.3%) 0.5750 (-10.1%) 25.0 (-8.1%)
TabM 0.8153 (-4.3%) 199.9 (-10.8%) 0.5760 (-11.3%) 318.9 (-20.3%)
TabM-PLE 0.8284 (-3.4%) 353.0 (-21.5%) 0.6085 (-7.4%) 212.2 (-38.8%)
Avg. -3.5% -8.7% -9.0% -18.9%
PFI MLP 0.8077 (-4.1%) 25.8 (+4.2%) 0.6262 (-1.2%) 17.2 (+4.8%)
MLP-PLE 0.8214 (-3.0%) 31.7 (+7.2%) 0.6393 (-0.9%) 27.2 (+6.9%)
TabM 0.8237 (-3.4%) 338.4 (-10.0%) 0.6392 (-1.8%) 415.0 (-6.7%)
TabM-PLE 0.8197 (-4.3%) 548.2 (-21.8%) 0.6451 (-1.8%) 299.6 (-21.7%)
Avg. -3.7% -5.1% -1.4% -4.2%
Summary. Table 3 highlights complementary strengths of the ShapG and PFI methods. [^2] Percentage changes in Table 3 are computed per-dataset and then averaged across all datasets. ShapG performs exceptionally well on classification tasks, achieving only an average accuracy reduction of 3.5% while delivering the largest reductions in training time across all models demonstrating its strong efficiency advantage. In contrast, the PFI method shows remarkable robustness in regression tasks, with an average drop of only 1.4% in $R^{2}$, making it highly reliable when predictive stability is essential. Taken together, these results show that the ShapG method offers superior computational benefits, whereas the PFI method excels in preserving regression performance, providing two complementary and effective strategies for feature selection.
4 Conclusions and future work
@@
In this work, we proposed a feature selection pipeline based on game-theoretic feature importance evaluation. The pipeline employs a $k$-means-based diversity sampling procedure to obtain a representative subset of observations, thereby reducing the influence of outliers and enabling more stable feature ranking. We evaluated two variants of this pipeline, namely, ShapG and PFI approaches, across ten benchmark datasets using state-of-the-art tabular deep-learning models. The results show that the proposed pipeline consistently reduces training cost while maintaining competitive predictive performance. As summarized in Table 3, ShapG offers the largest reduction in training time with only a modest loss in classification accuracy, whereas the permutation-based method exhibits higher robustness on regression tasks and causes only minimal degradation in the corresponding metrics. These observations indicate that ShapG is a suitable choice when computational efficiency is prioritized, while the permutation-based method is preferable when the stability of regression performance is of primary importance.
@@
Our study has several limitations that open opportunities for future research. First, we evaluated only two feature selection configurations (retaining the top 80% of features by importance) using the ShapG and PFI methods. Comprehensive comparisons with alternative feature selection approaches, including LIME and other game-theoretic methods with different characteristic functions, remain to be explored. We can also extend the proposed method by varying the number of features selected (the percentage of $q$-top features) for a fixed drop in the model performance. This extension is due to varying the number of significant features depending of the dataset and deep-learning model using as a basis.
@@
Additionally, extending this benchmark to industrial datasets, such as those described in rubachev2024tabred, would provide valuable insights on performance of real-world data analysis.
@@
The proposed diversity sampling method also shows promising results beyond feature importance evaluation. Preliminary experiments suggest its effectiveness for training data sampling, in particular, for domains with heterogeneous data distributions, e.g. wireless network data reported by base stations (measurement reports). While we have observed encouraging results on proprietary wireless communication datasets with TabM (gorishniy2024tabm) and FT-transformer (gorishniy2021revisiting) models, we cannot present these findings due to confidentiality constraints. Future work should conduct a systematic comparison between the proposed diversity sampling method and existing cluster-based approach (axiotis2024data), and explore these applications on publicly available datasets.
@@
These findings suggest that feature selection remains a valuable preprocessing step even for modern deep-learning approaches to tabular data, offering a practical way to data analysts to balance computational efficiency with respect to predictive performance.